Flounder Co. is building a new hockey arena at a cost of $2,460,000. It received a downpayment of $500,000 from local businesses to support the project, and now needs to borrow $1,960,000 to complete the project. It therefore decides to issue $1,960,000 of 10%, 10-year bonds. These bonds were issued on January 1, 2019, and pay interest annually on each January 1. The bonds yield 9%.

Prepare the journal entry to record the issuance of the bonds on January 1, 2019. (Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g. 58,971. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

January 1, 2019

Prepare a bond amortization schedule up to and including January 1, 2023, using the effective interest method. (Round answers to 0 decimal places, e.g. 38,548.)



Date


Cash
Paid


Interest
Expense


Premium
Amortization

Carrying
Amount of
Bonds

1/1/19 $

$

$

$

1/1/20

1/1/21

1/1/22

1/1/23

Assume that on July 1, 2022, Flounder Co. redeems half of the bonds at a cost of $1,057,800 plus accrued interest. Prepare the journal entry to record this redemption. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

July 1, 2022

(To record interest)

July 1, 2022

(To record reacquisition)
Answer #1

Journal Entry for issuance of bonds:

Date Account title and explanation Debit Credit
January 1,2019 Cash $2,085,785
Bonds payable $1,960,000
Premium on bonds payable $125,785
[To record issuance of bonds]

Calculations:

Interest payment = $1,960,000 x 10% = $196,000

Present value of the interest payments $1,257,861
[$196,000 x 6.41766 present value annuity factor (10%, 10 years)]
Present value of the face value $827,924
[$1,960,000 x 0.42241 present value factor (10%, 10 years)]
Cash received $2,085,785

Amortization Schedule:

Date Cash
Paid
Interest
Expense
Premium
Amortization
Carrying amount
of Bonds
1/1/19 $2,085,785
1/1/20 $196,000 $187,721 $8,279 $2,077,506
1/1/21 $196,000 $186,976 $9,024 $2,068,481
1/1/22 $196,000 $186,163 $9,837 $2,058,644
1/1/23 $196,000 $185,278 $10,722 $2,047,922

Cash paid = Interest payment

Interest expense = Preceding carrying amount x 9%

Premium amortization = Cash paid - Interest expense

Carrying amount = Preceding carrying amount - Premium amortization

Journal Entry to record redemption:

Date Account title and explanation Debit Credit
July 1,2022 Interest expense* $46,319
Premium on bonds payable $2,681
Cash** $49,000
[To record interest]
July 1,2022 Bonds payable (1,9600,000 x 1/2) $980,000
Premium on bonds payable (unamortized)*** $46,641
Loss on redemption of bonds $31,159
Cash $1,057,800
[To record reacquistion]

Calculations:

i.*Interest expense = ($2,058,644 x 1/2) x 9% x 6/12 = $1,029,322 x 9% x 6/12 = $92,639 x 6/12 = $46,319

ii.**Cash = (1,960,000 x 1/2) x 10% x 6/12 = $980,000 x 10% x 6/12 = $98,000 x 6/12 = $49,000

iii.***Unamortized premium = Carrying value of the bonds redeemed - Face value of the bonds redeemed

= [(2,058,644 x 1/2)-2,681] - [1,9600,000 x 1/2] = $1,032,003 - $980,000 = $46,641